OpenStudy (daniel.ohearn1):

Integrate ( (x^2 -2x -1) / (x-1)^2(x^2 +1) ) dx

OpenStudy (anonymous):

looks like a job for partial fractions, of the most annoying kind

OpenStudy (anonymous):

but it has been cooked up in this case so all your coefficients are easy

OpenStudy (daniel.ohearn1):

I found some values for A B C and D but idk if they are right they give me a shady answer..

OpenStudy (anonymous):

\[\frac{x^2-2x-1}{(x-1)^2(x^2+1)}\] \[=\frac{A}{x-1}+\frac{B}{(x-1)^2}+\frac{Cx+d}{(x^2+1}\] is a start

OpenStudy (anonymous):

did you get \[\frac{1}{x-1}+\frac{1}{(x-1)^2}+\frac{-x+1}{x^2+1}\]?

OpenStudy (anonymous):

when you say A,B,C, D i am not sure where you put them, so it is not clear

OpenStudy (anonymous):

if that is what you got, the integrals are easy enough

OpenStudy (daniel.ohearn1):

What are the values you got for the numerators?

OpenStudy (anonymous):

the ones i wrote above

OpenStudy (daniel.ohearn1):

A,B,C,D I mean

OpenStudy (anonymous):

like i said, it depends on how you label them the answer is above

OpenStudy (anonymous):

\[\frac{1}{x-1}+\frac{1}{(x-1)^2}+\frac{-x+1}{x^2+1}\]

OpenStudy (daniel.ohearn1):

It doesnt if you just tell which values you have numerically without mention of the alphabetical variables.

OpenStudy (anonymous):

three ones and a minus one which integral are you having trouble with?

OpenStudy (daniel.ohearn1):

I got different values but hold on Idk if I might get the right answer anyways...

OpenStudy (mathmale):

Keep in mind that you can choose any letters you want for the numerators of these partial fractions, and the order in which you use them doesn't matter. Note that the denominator of the integrand you were given factors out to (x-1)(x-1)(x^2+1). There are 3 factors here. You'll end up with 3 partial fractions. Note that (x-1)^2 is a "repeated root." That's why (x-1) and (x-1)^2 show up in two of the three partial fractions. Arbitrarily choosing letters for the numerators of these three partial fractions, we might get \[\frac{ P}{ x-1 }+\frac{ N }{ (x-1)^2 }+\frac{ Cx+F }{ x^2+1 }\]

OpenStudy (mathmale):

Need any further help with this problem? Have you been able to find the values of the four constant coefficients, whether A, B, C and D or P, N, C and F?

OpenStudy (daniel.ohearn1):

no not yet

OpenStudy (mathmale):

No, not yet (to finding the coeff. values), or no, not yet (to needing help)?

OpenStudy (daniel.ohearn1):

finding the coefficient values...

OpenStudy (mathmale):

OK. Have you a preference for the method used to find the values of those coefficients? \[\frac{ P}{ x-1 }+\frac{ N }{ (x-1)^2 }+\frac{ Cx+F }{ x^2+1 }=\frac{ x^2-2x-1 }{ (x-1)^2(x^2+1) }\]

OpenStudy (mathmale):

One (of several available methods) is this: Choose any four constants, such as 0, -1, 2, 3. One at a time, substitute said values for every occurrence of "x" into the equation given above. Note that you can't use x=1 for this purpose because you'd then have division by zero. Do this 3 more times, so you'll end up with four equations in four unknowns (P, N, C and F). Then this becomes a matrix problem where your job is to find the values of P, N, C and F. Making progress? or need any further assistance?

OpenStudy (mathmale):

If you were to let x=0, here's what your result would look like: \[\frac{ P}{ 0-1 }+\frac{ N }{ (0-1)^2 }+\frac{ Cx+F }{ 0^2+1 }=\frac{ 0^2-2(0)-1 }{ (0-1)^2(0^2+1) }\]

OpenStudy (mathmale):

Were you to let x=0, here's what you'd get: \[\frac{ P}{ 0-1 }+\frac{ N }{ (0-1)^2 }+\frac{ C(0)+F }{ 0^2+1 }=\frac{ 0^2-2(0)-1 }{ (0-1)^2(0^2+1) }\]

OpenStudy (mathmale):

You'd do this 3 more times, for x= -1, 2, 3.

OpenStudy (mathmale):

Still interested in finishing this problem? It's been 16 minutes since I last heard from you.

OpenStudy (daniel.ohearn1):

I got A=-1/4 B=-1 C=1/4 and D=-1/4 so we have A/x-1 + B/(x-1)^2 + Cx+D / (x^2 +1)

OpenStudy (mathmale):

You have come up with proposed numerical values for A, B, C and D. You need to substitute these numerical values for A, B, C and D in your expression on the far right: \[(-1/4)\frac{ 1 }{ x-1 }-1\frac{ 1 }{ (x-1)^2 }+\frac{ -1 }{ 4 }\frac{ x }{ x^2+1 }-\frac{ 1 }{ 4 }\frac{ x }{ x^2+1 }\]

OpenStudy (mathmale):

While this differs from the result satellite73 obtained and shared with you earlier, that doesn't mean you're not correct. Any questions or comments at this point?

OpenStudy (daniel.ohearn1):

I've got an answer but how did you get your answer if you didn't get these values?

OpenStudy (daniel.ohearn1):

And the LCD is (x^2 +1 )(x-1)?

OpenStudy (mathmale):

There are various methods in which to solve for the unknown coefficients. I don't know which method you used, nor do I know which method satellite73 used. My most recent post reflects YOUR work in that I substituted y our value for A into the expression with fractions, and so on. So, what are you asking me?

OpenStudy (mathmale):

I'm more interested in seeing / knowing that you know how to apply at least one of those methods for finding the coefficient values. Accuracy in applying it may come later.